A home security system is designed to have a 99% reliability rate. Suppose that nine home equipped with this system experience an attempted burglary. Find the probability that at least one of the alarms are triggered.

Group of answer choices

1- (.019 )

1- (.999 ) - 9* (.998 )* (.01)

1- (.999 )

1- (.019 ) - 9* (.018 )* (.99)
Answer #1

P(at least one alarms are triggered) = 1 - P(none of the nine alarms got triggered)

= 1 - (1 - 0.99)9

= 1 - (0.01)9

Option-A)